For the number of heads when 18 coins are tossed, find the following. Round your answers to three decimal places. Part 1 out of 2 Find the mean. Variance, and Standard deviation
Answer #1

Solution :

and q=(1-p) = 2 n=18, p==

Mean (μ)-n * p 18 * )-9

( :) * ( 1-1 ) -18 *-*--18 = 9= 4.5 1 1189 Variance (σ2)-n*p*q= n*p*(1-p)-18 * Variance (σ2) = n*p*q = n*p*(1-p) = 18* ) * (1

St )-Va2_ V4.5-212132~2.121 andard deviation σ